# MnLargeSymbols’164 MnLargeSymbols’171

### v1

@@ Abstract
The choice of batch sizes in stochastic gradient optimizers is critical for model training. However, the practice of varying batch sizes throughout the training process is less explored compared to other hyperparameters. We investigate adaptive batch size strategies derived from adaptive sampling methods, traditionally applied only in stochastic gradient descent. Given the significant interplay between learning rates and batch sizes, and considering the prevalence of adaptive gradient methods in deep learning, we emphasize the need for adaptive batch size strategies in these contexts. We introduce AdAdaGrad and its scalar variant AdAdaGrad-Norm, which incrementally increase batch sizes during training, while model updates are performed using AdaGrad and AdaGrad-Norm. We prove that AdaGrad-Norm converges with high probability at a rate of $\mathscr{O}(1/K)$ for finding a first-order stationary point of smooth nonconvex functions within $K$ iterations. AdaGrad also demonstrates similar convergence properties when integrated with a novel coordinate-wise variant of our adaptive batch size strategies. Our theoretical claims are supported by numerical experiments on various image classification tasks, highlighting the enhanced adaptability of progressive batching protocols in deep learning and the potential of such adaptive batch size strategies with adaptive gradient optimizers in large-scale model training.
@@ 1 Introduction
Large-scale optimization algorithms (Bottou et al., 2018) form the foundation of the success of deep learning in the generative AI era. Stochastic gradient descent (SGD; Robbins and Monro, 1951) and its many variants, along with batch sampling techniques, are pivotal in training deep learning models. However, training deep neural networks, such as those in transformers, is notoriously challenging due to their high dimensionality and nonconvex landscape. This complexity necessitates extensive hyperparameter tuning and sophisticated training strategies for the optimizers to circumvent premature divergence and training instabilities. Consequently, training deep learning models often appears more as an art than a science. The most critical hyperparameter is arguably the learning rate (or step size). Adaptive gradient methods with adaptive learning rates, like AdaDelta (Zeiler, 2012), AdaGrad (Duchi et al., 2011), and Adam (Kingma and Ba, 2015), are now prevalent, as they reduce the need for meticulous tuning and complex learning rate schedules typically required for SGD. Another important, yet frequently overlooked hyperparameter is the batch size. It moderates the trade-off between computational efficiency and model generalization by controlling the stochastic noise in batch gradients. Nonetheless, the selection of batch sizes in deep learning training largely remains heuristic, such as using a constant size for convolutional networks or a linear warm-up for large language models (Brown et al., 2020; Rae et al., 2021; Hoffmann et al., 2022), and is usually predetermined before training begins. Additionally, the intricate relationship between learning rates and batch sizes warrants attention. Specifically, Smith et al. (2018) demonstrated the equivalence of reducing step sizes and increasing batch sizes, but this principle mainly applies to SGD. The impact of varying batch sizes in adaptive gradient methods has not yet been fully explored.
In light of this, our work seeks to unravel the complexities surrounding the determination of optimal batch sizes for adaptive gradient methods. We aim to introduce automated schedules capable of dynamically deciding when to increase the batch size and by how much, based on the training needs. Our approach is theoretically grounded and relies on the statistics of batch gradients of iterates, thereby adapting to the training dynamics of various model types.
@@
Our proposed strategies are based on adaptive sampling methods (Byrd et al., 2012; Bollapragada et al., 2018a). In the context of deep learning, De et al. (2016, 2017) have numerically shown the effectiveness of these methods with AdaDelta. However, the convergence properties of such adaptive sampling methods for adaptive gradient methods have not been thoroughly investigated, leaving a gap between theory and practice. Moreover, in existing adaptive sampling methods developed mainly for SGD, learning rates are often fixed or adjusted using backtracking line search procedures (mainly for convex problems). This becomes computationally impractical or inefficient for larger models, especially given the nonconvex nature of deep neural networks.
@@
The development of adaptive batch size strategies for deep learning is not a novel concept, featuring methodologies like Big Batch SGD (De et al., 2016, 2017), CABS (Balles et al., 2017), AdaBatch (Devarakonda et al., 2017), SimiGrad (Qin et al., 2021), and AdaScale SGD (Johnson et al., 2020), which adapts learning rates for large batches rather than directly adjusting batch sizes. However, these methods often lack a principled basis with rigorous convergence guarantees or are limited to analyses of SGD under stringent conditions (e.g., convexity or the Polyak–Łojasiewicz condition), despite the prevalence of adaptive gradient methods in deep learning. Furthermore, these approaches still necessitate the choice of learning rate adjustments, such as line search routines or schedulers, leaving a gap in achieving full adaptivity. Additionally, their strategies for determining new batch sizes may rely on heuristic decisions, such as geometric growth or decay rates as suggested by Qin et al. (2021).
@@
In this work, we aim to bridge the gaps in the theory and practice of adaptive batch size schemes for machine learning. We introduce two adaptive batch size schemes derived in a principled manner, grounded in the so-called adaptive sampling methods (Byrd et al., 2012), tailored for adaptive gradient methods. Our focus is on AdaGrad and its norm variant, AdaGrad-Norm, which are among the simplest and most extensively studied adaptive gradient methods. The development of adaptive batch size schemes for these methods holds significant theoretical and practical relevance.
From a theoretical perspective, we establish a sublinear convergence rate (with high probability) for our proposed methods when applied with AdaGrad-Norm and AdaGrad to smooth nonconvex objectives. This significantly broadens the existing body of work, which has primarily focused on SGD. Moreover, we relax the Lipschitz smoothness condition on the objective function by adopting the generalized smoothness concept introduced by Zhang et al. (2020b, a). This adaptation allows for more generalized and realistic applications in the current practice of deep learning. On the empirical front, we demonstrate the efficacy of our proposed methods through various small-scale experiments on image classification tasks. These experiments highlight the benefits of our schemes’ adaptivity, driven by both adaptive batch sizes and learning rates.
Organization.
We delve into related work in Section 2. In Section 3, we define the problem of interest and provide an overview of adaptive gradient methods. Following this, we detail the adaptive sampling methods in Section 4 and establish the convergence of AdAdaGrad-Norm and AdAdaGrad in Section 5. Small-scale numerical experiments are presented in Section 6. We conclude our paper by summarizing our contributions and discussing potential extensions in Section 7.
@@ 2 Related Work
We give an overview of related work on batch sizes in model training and adaptive sampling methods. Further related work can be found in Appendix D.
2.1 Effect of Batch Sizes on Model Training
@@
In stochastic gradient optimizers, batch sizes play a crucial role in controlling the variance of stochastic gradients as estimators of the full deterministic gradients. Although the noise from stochastic gradients offers benefits for nonconvex optimization, the trend towards using larger batches in large-scale model training has become standard, thanks to advancements in hardware, significantly reducing training time. The concept of large batch training in deep learning was popularized by the work of Smith and Le (2018); Smith et al. (2018) and has been widely adopted in fields such as computer vision (Goyal et al., 2017) and natural language processing (You et al., 2020). The impact of batch sizes has been further examined in a more principled manner in subsequent studies. For example, McCandlish et al. (2018) introduced an empirical model for large batch training without providing rigorous proof, postulating the existence of critical batch sizes through extensive numerical simulations on convolutional neural networks (CNNs), LSTMs, and VAEs. Meanwhile, Kaplan et al. (2020) focused their experiments on transformers. Zhang et al. (2019) explored how critical batch sizes vary with the nature of optimizers, including acceleration (momentum), preconditioning, and exponential moving average, through both large-scale experiments and a noisy quadratic model. Granziol et al. (2022) applied random matrix theory to examine the batch Hessian, theoretically establishing learning rates as a function of batch size (i.e., the scaling rule) for both SGD (linear) and adaptive gradient methods (square root), with experimental validation. Furthermore, Shallue et al. (2019) investigated the impact of batch sizes in the context of data parallelism and empirically characterized the effects of large batch training, finding no evidence of a degradation in generalization performance.
2.2 Adaptive Sampling Methods
@@
In the realm of stochastic optimization, a more theoretically grounded approach known as adaptive sampling methods has been developed for batch (or minibatch) algorithms. Byrd et al. (2012) introduced a method termed the norm test, which adaptively increases the batch size throughout the optimization process. The rationale behind the norm test traces back to Carter (1991) and represents a more generalized condition compared to the approach by Friedlander and Schmidt (2012). Both approaches demonstrate linear convergence when the batch size increases geometrically. Bollapragada et al. (2018a) proposed the augmented inner product test, which allows for more gradual increases in batch sizes than the norm test. Additionally, Cartis and Scheinberg (2018) introduced a relaxation of the norm test that permits its condition to be violated with a probability of less than $0.5$. This suite of adaptive sampling methods is considered part of the variance-reduced optimization algorithms (Johnson and Zhang, 2013; Reddi et al., 2016) that are widely used in machine learning (Gower et al., 2020).
Adaptive sampling methods have also been extended to tackle problems outside the unconstrained optimization setting. Xie et al. (2023) proposed a proximal extension of the norm test and the inner product test to minimize a convex composite objective, which comprises a stochastic function and a deterministic, potentially nonsmooth function. Beiser et al. (2023) explored deterministic constrained problems, including those with nonconvex objectives. Moreover, adaptive sampling methods have been applied to various optimization problems and algorithms, such as Bollapragada et al. (2018b) for L-BFGS, Berahas et al. (2022) for sequential quadratic programming (SQP) in the context of equality-constrained problems, Bollapragada et al. (2023) for augmented Lagrangian methods, and Bollapragada and Wild (2023) for quasi-Newton methods.
3 Preliminaries
@@
3.1 Problem Formulation
We consider the problem of minimizing the expected risk $\mathbb{E}_{\xi\sim\mathcal{P}}[f(x;\xi)]$ with respect to $x\in\mathbb{R}^{d}$, where the random variable $\xi\in\mathcal{Z}\subseteq\mathbb{R}^{p}$ is distributed according to the probability distribution $\mathcal{P}$. Since the distribution $\mathcal{P}$ is unknown, we approximate it with the empirical distribution $\widehat{\mathcal{P}}=\frac{1}{n}\sum_{i=1}^{n}\updelta_{\xi_{i}}$, where $\{\xi_{i}\}_{i\in\llbracket n\rrbracket}$ is the sample of size $n\in\mathbb{N}^{*}$. This leads to the following empirical risk minimization problem:
$\operatorname*{minimize}_{x\in\mathbb{R}^{d}}\ F(x)\coloneqq\mathbb{E}_{\xi\sim\widehat{\mathcal{P}}}[f(x;\xi)]=\frac{1}{n}\sum_{i=1}^{n}f(x;\xi_{i}).$
In practice, the sample size $n$ is usually very large and the loss function $F$ and its gradient are approximated by their batch counterparts. Given a batch of samples $\mathcal{B}\subset\llbracket n\rrbracket$ of size $b\coloneqq|\mathcal{B}|$, we define the batch loss associated with the batch $\mathcal{B}$ by $F_{\mathcal{B}}(x)\coloneqq\frac{1}{b}\sum_{i\in\mathcal{B}}f(x;\xi_{i})$. If $f(\cdot;\xi)$ is continuously differentiable for any $\xi\in\mathcal{Z}$, the gradients of the loss and the batch loss are, respectively, given by
$\nabla F(x)=\frac{1}{n}\sum_{i=1}^{n}\nabla f(x;\xi_{i}),\;\nabla F_{\mathcal{B}}(x)=\frac{1}{b}\sum_{i\in\mathcal{B}}\nabla f(x;\xi_{i}).$
The batch gradient is an unbiased estimator for the loss function gradient, $\mathbb{E}_{\mathcal{B}}[\nabla F_{\mathcal{B}}(x)]=\nabla F(x)$ for any $x\in\mathbb{R}^{d}$. In many applications, including deep learning, the objective function $F$ is nonconvex. Thus, we consider the problem of finding a (first-order) $\varepsilon$-stationary point $\widetilde{x}\in\mathbb{R}^{d}$ that satisfies $\|\nabla F(\widetilde{x})\|^{2}\leqslant\varepsilon$, rather than a global minimum that is intractable.
3.2 Adaptive Gradient Methods
We focus on two simple adaptive gradient methods, AdaGrad (Duchi et al., 2011; McMahan and Streeter, 2010) and AdaGrad-Norm (Streeter and McMahan, 2010), whose step sizes are computed adaptively based on the magnitudes of previous (batch) stochastic gradients.
AdaGrad was proposed for online convex optimization, which takes the following update:
$(\forall k\in\mathbb{N}^{*})\quad\begin{aligned} v_{k}&=v_{k-1}+g_{k}^{2}\\ x_{k+1}&=x_{k}-\alpha g_{k}\odot v_{k}^{-\nicefrac{{1}}{{2}}},\end{aligned}$ (1)
where $g_{k}\coloneqq\nabla F_{\mathcal{B}_{k}}(x_{k})$, $\alpha\in\mathbb{R}_{++}$ is a constant step size (or learning rate), $\odot$ denotes the Hadamard product, and the power operations are performed coordinate-wise. Since $(v_{k})_{k\in\mathbb{N}}\subset\mathbb{R}_{++}^{d}$ is a sequence of positive vectors, AdaGrad has adaptive coordinate-wise step sizes.
$(\forall k\in\mathbb{N}^{*})\quad\begin{aligned} v_{k}&=v_{k-1}+\left\lVert g_{k}\right\rVert^{2}\\ x_{k+1}&=x_{k}-\alpha g_{k}/\sqrt{v_{k}},\end{aligned}$ (2)

### v2

@@ Abstract
The choice of batch sizes in minibatch stochastic gradient optimizers is critical in large-scale model training for both optimization and generalization performance. Although large-batch training is arguably the dominant training paradigm for large-scale deep learning due to hardware advances, the generalization performance of the model deteriorates compared to small-batch training, leading to the so-called “generalization gap” phenomenon. To mitigate this, we investigate adaptive batch size strategies derived from adaptive sampling methods, originally developed only for stochastic gradient descent. Given the significant interplay between learning rates and batch sizes, and considering the prevalence of adaptive gradient methods in deep learning, we emphasize the need for adaptive batch size strategies in these contexts. We introduce AdAdaGrad and its scalar variant AdAdaGrad-Norm, which progressively increase batch sizes during training, while model updates are performed using AdaGrad and AdaGrad-Norm. We prove that AdAdaGrad-Norm converges with high probability at a rate of $\mathscr{O}(1/K)$ to find a first-order stationary point of smooth nonconvex functions within $K$ iterations. AdAdaGrad also demonstrates similar convergence properties when integrated with a novel coordinate-wise variant of our adaptive batch size strategies. We corroborate our theoretical claims by performing image classification experiments, highlighting the merits of the proposed schemes in terms of both training efficiency and model generalization. Our work unveils the potential of adaptive batch size strategies for adaptive gradient optimizers in large-scale model training.
@@ 1 Introduction
Large-scale optimization algorithms (Bottou et al., 2018) form the foundation of the success of deep learning in the era of generative AI. Minibatch stochastic gradient descent (SGD) (Robbins and Monro, 1951) and its many variants, along with batch sampling techniques, are the main workhorse for training deep neural networks. However, training deep neural networks, such as those in transformers, is notoriously challenging because of their high dimensionality and nonconvex landscape. This complexity necessitates extensive hyperparameter tuning and sophisticated training strategies for optimizers to circumvent premature divergence and training instabilities. Consequently, training deep learning models often appears as more an art than a science. The most critical hyperparameter is arguably the learning rate (or step size). Adaptive gradient methods with adaptive learning rates, such as AdaDelta (Zeiler, 2012), AdaGrad (Duchi et al., 2011), and Adam (Kingma and Ba, 2015), are now prevalent, as they reduce the need for meticulous tuning and complex learning rate schedules, typically required for SGD. Another important, yet frequently overlooked hyperparameter is the batch size. It moderates the trade-off between computational efficiency and model generalization by controlling the magnitude of noise in batch gradients. However, the selection of batch sizes in deep learning training remains largely heuristic, such as using a constant size for convolutional networks or a linear warm-up for large language models (Brown et al., 2020; Rae et al., 2021; Hoffmann et al., 2022), and is usually predetermined before training begins. Furthermore, from a hardware utilization perspective, the use of a large number of distributed computational resources (i.e., GPUs or TPUs) necessitates the large-batch training regime when parallel minibatch SGD (Zinkevich et al., 2010; Dean et al., 2012) is used. In addition, the intricate relationship between learning rates and batch sizes deserves attention. Specifically, Smith et al. (2018) demonstrated the equivalence of reducing step sizes and increasing batch sizes, but this principle applies mainly to SGD. The impact of varying batch sizes in adaptive gradient methods has not yet been fully explored.
@@
In light of this, our work seeks to unravel the complexities surrounding the determination of optimal batch sizes for adaptive gradient methods. We aim to introduce automated schedules capable of dynamically deciding when to increase the batch size and by how much, based on the training needs. Our approach is theoretically grounded and relies on the statistics of batch gradients of iterates, thereby adapting to the training dynamics of the models. Our proposed schedules are able to reduce the generalization gap in large-batch training, but at the same time also maintain the utilization of large batches at the later stage of the overall training process. Our proposed strategies are based on adaptive sampling methods (Byrd et al., 2012; Bollapragada et al., 2018a). In the context of deep learning, De et al. (2016, 2017) have numerically shown the effectiveness of these methods with AdaDelta. However, the convergence properties of such adaptive sampling methods for adaptive gradient methods have not been thoroughly investigated, leaving a gap between theory and practice. Moreover, in existing adaptive sampling methods developed mainly for SGD, step sizes are often fixed or adjusted using backtracking line search procedures (mainly for convex problems). This becomes computationally impractical or inefficient for larger models, especially given the nonconvex nature of deep neural networks.
@@
The development of adaptive batch size strategies for deep learning is not a novel concept, featuring methodologies such as Big Batch SGD (De et al., 2016, 2017), CABS (Balles et al., 2017), AdaBatch (Devarakonda et al., 2017), SimiGrad (Qin et al., 2021), and AdaScale SGD (Johnson et al., 2020), which adapts learning rates for large batches rather than directly adjusting batch sizes. However, these methods often lack a principled basis with rigorous convergence guarantees or are limited to analyses of SGD under stringent conditions (e.g., convexity or the Polyak–Łojasiewicz condition), despite the prevalence of adaptive gradient methods in deep learning. Furthermore, these approaches still necessitate the choice of learning rate adjustments, such as line search routines or schedulers, leaving a gap in achieving full adaptivity. Furthermore, strategies for determining new batch sizes may rely on heuristic rules, such as geometric growth or decay rates (Qin et al., 2021).
@@
In this work, our objective is to bridge the theoretical and practice gaps of adaptive batch size schemes for adaptive gradient methods. We introduce two principal adaptive batch size schemes, grounded in the so-called adaptive sampling methods (Byrd et al., 2012; Friedlander and Schmidt, 2012; Bollapragada et al., 2018a), tailored for adaptive gradient methods. Our focus is on AdaGrad and its norm variant AdaGrad-Norm, which are among the simplest and most extensively studied adaptive gradient methods. The development of adaptive batch size schemes for these methods has significant theoretical and practical relevance.
The technical contributions of this work are three-fold. From a theoretical perspective, we establish a sublinear convergence rate (with high probability) for our proposed methods when applied with AdaGrad-Norm and AdaGrad to smooth nonconvex objectives. This significantly broadens the existing body of work which has mainly focused on SGD. Moreover, we relax the Lipschitz smoothness condition on the objective function by adopting the generalized smoothness concept introduced in Zhang et al. (2020b, a). This adaptation allows for more generalized and realistic applications in the current practice of deep learning. On the empirical front, we demonstrate the efficacy of our proposed methods through various numerical experiments on image classification tasks. These experiments highlight the benefits of the adaptivity of our schemes, driven by both adaptive batch sizes and step sizes. Finally, we provide an efficient implementation of the proposed approach in PyTorch, making use of the torch.func module for efficient parallelized computations of per-sample gradients using the so-called vectorizing map function vmap. An open source code repository will be provided upon acceptance. To the best of our knowledge, our proposed methods are the first adaptive batch size schemes based on adaptive sampling methods for adaptive gradient methods, which are convergent with guarantees, empirically performant and efficiently implemented in deep learning libraries.
@@ 2 Related Work
We give an overview of related work on batch sizes in model training and adaptive sampling methods. Further related work can be found in Appendix A.
Large-batch training.
@@
In stochastic gradient optimizers, batch sizes play a crucial role in controlling the variance of stochastic gradients as estimators of the full deterministic gradients. Although the noise from stochastic gradients offers benefits for nonconvex optimization, the trend towards using larger batches in large-scale model training has become standard, thanks to advancements in parallelized hardware such as GPUs, significantly reducing training time. The concept of large-batch training in deep learning was popularized in Smith and Le (2018); Smith et al. (2018) and has been widely adopted in applications such as ImageNet classification (Goyal et al., 2017) and BERT training (You et al., 2020). In the work of LeCun et al. (2002), it has been understood that model generalization deteriorates in large-batch training. Large-batch training tends to result in a loss landscape with many sharp minima, which are harder to escape during optimization and hence lead to worse generalization (Keskar et al., 2017). The impact of batch sizes has been further examined in a more principled manner in subsequent studies. For example, McCandlish et al. (2018) introduced an empirical model for large batch training without providing rigorous proof, postulating the existence of critical batch sizes through extensive numerical simulations on convolutional neural networks (CNNs), LSTMs, and VAEs. Meanwhile, Kaplan et al. (2020) focused their experiments on transformers. Zhang et al. (2019) explored how critical batch sizes vary with the nature of optimizers, including acceleration (momentum), preconditioning, and exponential moving average, through both large-scale experiments and a noisy quadratic model. Granziol et al. (2022) applied random matrix theory to examine the batch Hessian, theoretically establishing learning rates as a function of batch size (i.e., the scaling rule) for both SGD (linear) and adaptive gradient methods (square root), with experimental validation. Although the “generalization gap” can be narrowed by using larger learning rates proportional to batch sizes to maintain the gradient noise scale (Hoffer et al., 2017; Smith and Le, 2018; Smith et al., 2018), it cannot be completely eliminated. However, Shallue et al. (2019) investigated the impact of batch sizes in the context of data parallelism and empirically characterized the effects of large-batch training, finding no evidence of a degradation in generalization performance.
Adaptive sampling methods.
@@
In the realm of stochastic optimization, a more theoretically grounded approach known as adaptive sampling methods has been developed for batch (or minibatch) algorithms. Byrd et al. (2012) introduced a method termed the norm test, which adaptively increases the batch size throughout the optimization process. The rationale behind the norm test traces back to Carter (1991) and represents a more generalized condition compared to the approach of Friedlander and Schmidt (2012). Both approaches demonstrate linear convergence when the batch size increases geometrically. Bollapragada et al. (2018a) proposed the augmented inner product test, which allows for more gradual increases in batch sizes than the norm test. Furthermore, Cartis and Scheinberg (2018) introduced a relaxation of the norm test that allows its condition to be violated with a probability of less than $0.5$. This suite of adaptive sampling methods is considered part of the variance-reduced optimization algorithms (Johnson and Zhang, 2013; Reddi et al., 2016) that are widely used in machine learning (Gower et al., 2020).
3 Problem Formulation
@@
Problem setting.
We consider the problem of minimizing the expected risk $\mathbb{E}_{\xi\sim\mathsf{P}}[f(x;\xi)]$ with respect to $x\in\mathbb{R}^{d}$, where the random variable $\xi\in\mathcal{Z}\subseteq\mathbb{R}^{p}$ is distributed according to the unknown true data distribution $\mathsf{P}$. We approximate $\mathsf{P}$ with the empirical distribution $\widehat{\mathsf{P}}=\frac{1}{n}\sum_{i=1}^{n}\updelta_{\xi_{i}}$, where $\{\xi_{i}\}_{i\in\llbracket n\rrbracket}$ is the sample of size $n\in\mathbb{N}^{*}$, which leads to the empirical risk minimization problem:
$\operatorname*{minimize}_{x\in\mathbb{R}^{d}}\ F(x)\coloneqq\mathbb{E}_{\xi\sim\widehat{\mathsf{P}}}[f(x;\xi)]=\frac{1}{n}\sum_{i=1}^{n}f(x;\xi_{i}).$
When the sample size $n$ is large, the gradient of $F$ is approximated by its batch counterparts. Given a batch of samples $\mathcal{B}\subset\llbracket n\rrbracket$ of size $b\coloneqq|\mathcal{B}|$, we define the batch loss associated with the batch $\mathcal{B}$ by $F_{\mathcal{B}}(x)\coloneqq\frac{1}{b}\sum_{i\in\mathcal{B}}f(x;\xi_{i})$. If $f(\cdot;\xi)$ is continuously differentiable for any $\xi\in\mathcal{Z}$, the loss and batch loss gradients are given, respectively, by $\nabla F(x)=\frac{1}{n}\sum_{i=1}^{n}\nabla f(x;\xi_{i})$ and $\nabla F_{\mathcal{B}}(x)=\frac{1}{b}\sum_{i\in\mathcal{B}}\nabla f(x;\xi_{i})$. The batch gradient is an unbiased estimator for the loss function gradient, i.e., $\mathbb{E}_{\mathcal{B}}[\nabla F_{\mathcal{B}}(x)]=\nabla F(x)$ for any $x\in\mathbb{R}^{d}$. In many applications, including deep learning, the objective function $F$ is nonconvex. Thus, we consider the problem of finding a (first-order) $\varepsilon$-stationary point $x^{\star}\in\mathbb{R}^{d}$ that satisfies $\|\nabla F(x^{\star})\|^{2}\leqslant\varepsilon$, rather than a global minimum that is intractable.